OpenStudy (anonymous):

what is the anti derivative of 3x^2 and 4x?

OpenStudy (anonymous):

use the power rule backwards \[\int x^n=\frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

\[\int 3x^2dx = 3\frac{x^{2+1}}{2+1}=3\frac{x^3}{3}=x^3\]

OpenStudy (anonymous):

i guess you are supposed to put \(+C\) at the end

OpenStudy (anonymous):

guest celebrity on the scene!!!